OpenStudy (anonymous):

In how many different ways can you answer a multiple-choice test that has 5 questions and 3 choices for each answer? a)15 b)125 c)243 ithink it's a, i just need to make sure

hartnn (hartnn):

yes it is 15.....

OpenStudy (anonymous):

k, thanks

ganeshie8 (ganeshie8):

it should be 3*3*3*3*3 = 3^5 = 243 tricky one ! :D

hartnn (hartnn):

how? is it of the type {AAAAA,AAAAB......CCCCC} with number of elements in that set? then will it be 243?